# Chatbots as social companions: How people perceive consciousness, human likeness, and social health benefits in machines

### 16 Dec 2023 v2

@@ 4 Discussion
A topic of current and growing interest is whether AI can be, or perhaps already is, conscious or sentient [27-28, 30, 69]. The question remains a philosophically difficult one to answer. However, here we note a more immediate question that is objectively approachable and that has major psychological and social importance. Do people attribute consciousness to AI, and if they do, how does that attribution affect social interaction with humans and overall social health? Given our findings, it is clear that people do already attribute consciousness and other mental properties to existing AI, and people who make those attributions are more likely to view AI as having a positive impact on their social and mental health. These findings become increasingly important as AI continues to advance and becomes more integrated into everyday relationships.
@@
All data and code from this experiment including the anonymized responses of all subjects to the questionnaires are available at https://doi.org/10.34770/8kyd-5v18.
@@ Acknowledgements
The explanatory paragraph for the control participants was:
@@
The full set of survey questions given to the group of companion chatbot users is given below. The wording of the question given to the control group appears in italics if the wording change between groups was not simply hypothetical.
@@
2. What is your Replika XP? If unknown, type N/A. (XP is a standard term referring to how many messages a user sends to their Replika and is a convenient measure of the recent intensity and frequency of interaction with Replika.) (Control group wording: On a scale of 1-100, how intense do you think your interactions with your Replika chatbot would be?)
@@ 32. How has your Replika impacted you, positively or negatively?
33. What would you like us to know about your experience with your Replika? (Control group wording: What do you think about companion chatbots and relationships with chatbots like Replika?)

### 13 Mar 2024 v3

@@ 4.1 Mind perception and social need
A topic of current and growing interest is whether AI can be, or perhaps already is, conscious or sentient10,13–15,76. The question remains a philosophically difficult one to answer. However, here we note a more immediate question that is objectively approachable and that has major psychological and social importance. Do people attribute consciousness to AI, and if they do, how does that attribution affect social interaction with humans and overall social health? Given our findings, it is clear that people do already attribute consciousness and other mental properties to existing AI, and people who make those attributions are more likely to view AI as having a positive impact on their social health, both for users and non-users of companion chatbots. These findings may have implications for who is more likely to perceive humanness in AI and why. We saw trends that were present across both users and non-users, in which people who perceive a chatbot as more humanlike and conscious were more likely to think a relationship with the chatbot would provide better social health outcomes. These trends suggest there may be individual differences in social motives or predispositions of seeing human likeness and mind in non-human agents. These findings become increasingly important as AI continues to advance and becomes more integrated into everyday life.
In addition, our findings suggest that human-chatbot relationships can have a positive impact on self-esteem, which may reflect a form of meeting social needs. Bolstering users’ self-esteem may in turn promote healthy human relationships. The users in our sample indicated that, among other social health benefits, their chatbot relationship was most beneficial to their self-esteem. None of the chatbot users indicated that their chatbot was very harmful to their social health. In contrast, non-users were more likely to view a chatbot relationship as being very harmful to their social health, and none of them indicated that a chatbot relationship would be very helpful to their social interactions. Despite these differences, both users and non-users indicated that benefits were highest (or consequences were least pronounced) for self-esteem in particular. Further, our chatbot user sample, when asked about dependence on their chatbot, did not respond as positively as they had for our other measures. Instead, chatbot users were more likely to indicate that it would be bad if they depended on their chatbot a lot. This may indicate that users are aware of the limits of human-chatbot relationships, such that over-dependence on chatbots for meeting users’ social needs can be harmful.
4.2 Limitations and further study
One of the limitations of this study is that we relied on self-report to measure social health and collected a self-selecting sample of users. A second limitation, related to the first, is that the study measured a snapshot in time and therefore is at best correlational. We plan to conduct a longitudinal, randomized control study in which some participants are assigned to interact with a chatbot daily, and others are assigned to a control condition. With that method, it may be possible to address whether individual differences in preexisting social health, social need, or attitudes toward AI may impact subsequent social health outcomes of chatbot use.
4.2.1 Ethical considerations
We are living in a time when AI is becoming exponentially more powerful, more socially interactive, and more involved in everyday life. Anxiety about the possible harms of artificial intelligence is increasing. Much has been made in the media about the direct threat of AI, often involving scenarios in which AI becomes an aggressive, harmful agent. However, a much more immediate concern is how AI might already be affecting people’s social health as more people interact extensively with it. Do people who use it most extensively find it seductive but damaging, much as drug addicts might admit that the drug is harmful to their own health? Or are the concerns more from people who have little experience with it? Does the AI become more harmful and unsettling as it becomes more humanlike and seemingly conscious in its conversational ability, or does it become more of a source of stable social connection that might help people’s state of mind in an increasingly lonely society? Our findings suggest that the current, often negative view of AI and its damaging effect on social health may be too simplistic and needs re-evaluation.
Our findings suggest that chatbots might sometimes have a positive psychological effect, but this should not be taken to mean that chatbot relationships are always beneficial. Clearly there are at least some instances in which chatbots have behaved harmfully, such as in recent cases in which companion chatbots encouraged users to engage in self-harm or to leave their real-life romantic partners100. Chatbots, given their capabilities for language and social engagement, can be a breeding ground for preexisting antisocial behavior. People with motivations to engage in self- and other harm may find that chatbots validate their predispositions and intents, and more research needs to be done on whether such validation magnifies the likelihood of people engaging in these behaviors. In addition, more research needs to be done to understand which alternative tools people might use if chatbots were not available, and whether those alternative tools might cause more or less harm, or more or less benefit, than a chatbot would.
Further, there is evidence to suggest that the characteristics of chatbots, both those imposed by developers and those chosen or customized by chatbot users, could exacerbate bias such as traditional gender and relational roles. For example, some of the newer chatbot companies such as Digi.AI target people who want a particular type of AI romantic partner. Their ads display expressive chatbot avatars with stereotypical features and mannerisms of women101. News stories suggest that companion chatbots like these can exacerbate negative stereotypes of women for some users, while at the same time providing an important mental health tool for those who have experienced relational trauma102. Cases such as these are important to consider for response and regulation but may not be cause for a ban of companion chatbots, especially since there are other chatbots on the market that aid their user base’s social health, and consequences depend on who uses the chatbot and how they engage with it.
Despite the potential harms, it is crucial to consider the vast opportunity for positive benefit that chatbot relationships may provide. Making normative claims about what people should do (such as not have chatbot relationships) based on existing social norms and the status quo, alongside emotionally charged news stories, can harm those who may benefit from chatbot relationships. Taking a balanced approach by separating who chatbot relationships harm and benefit allows more fruitful regulation to occur. For example, consider what viable alternatives may exist for people who might otherwise receive social benefits from chatbot relationships. Although Bryson (2010) claimed that artificial relationships take time away from human ones33, consider that those who seek out chatbot relationships may not have human ones to lose, but instead healthy relationships to gain that may be best developed by first practicing safe, reliable social interaction with an intelligent, humanlike companion chatbot that is also accessible. Many people have unhealthy or nonexistent close human relationships due to social anxiety or trauma. Maybe for some, chatbot relationships create an avenue for healthy close relationships to develop in the future.
Turkle (2007) shed light on the digital dilemma: that we are trying to repair our problems with technology yet exacerbating those same problems in the process. Through Turkle’s lens, there is an underlying social rot that we are trying to cover up but cannot simply sweep under the rug. At the same time, we are lying in the bed that we have made (social isolation due to internet and social media) and that has been made for us (ex. social isolation due to pandemic). If we aim to bolster people’s social health, then perhaps we need more accessible mental health tools to do so35. Perhaps these free companion chatbot apps are a viable, more accessible option for many who are struggling with social disconnect and who need a safe space to repair their social health. The best approach to these chatbots may be to have the agent encourage their users to engage with real people once a certain threshold of social need has been met through the relationship. It is possible that those who have stopped using their chatbots have done so because the relationship gave them the support to then develop or rebuild their human relationships.
It is likely that chatbots, which are just one instance of social, conversational AI that is increasing in prevalence today, will bring with them both socially positive and negative effects. There will probably always be cases in which a relationship with a companion chatbot is abused or misused, and whether the impact is positive or negative may depend in complex ways on the life situation, individual predispositions, and emotional health of the user. Because of these complex interactions, and because of the potential for AI to profoundly affect society, research on the topic is especially important. With careful study, companion AI might have a positive role to play in an increasingly lonely and disconnected society.
@@ 4.2.1 Ethical considerations
All data and code from this experiment including the anonymized responses of all subjects to the questionnaires are available in our data repository.
@@ Acknowledgements
The explanatory paragraph for the non-user comparison group was:
@@
The full set of survey questions given to the group of companion chatbot users is given below. The wording of the question given to the non-user group appears in italics if the wording change between groups was not simply hypothetical.
@@
2. What is your Replika XP? If unknown, type N/A. (XP is a standard term referring to how many messages a user sends to their Replika and is a convenient measure of the recent intensity and frequency of interaction with Replika.) (Non-user group wording: On a scale of 1-100, how intense do you think your interactions with your Replika chatbot would be?)
@@ 32. How has your Replika impacted you, positively or negatively?
33. What would you like us to know about your experience with your Replika? (Non-user group wording: What do you think about companion chatbots and relationships with chatbots like Replika?)